# Learning Disease State from Noisy Ordinal Disease Progression Labels

### v1

@@ 3.1 Disease Progression Classification
Table 1: Classification performance for disease progression on the MARIO dataset.
@@
naïve classifier $70\pm 2$ $44\pm 2$ $87\pm 0$ $60\pm 1$ $57\pm 1$ $60\pm 1$
ours $61\pm 3$ $46\pm 0$ $88\pm 0$ $59\pm 2$ $57\pm 0$ $62\pm 0$
ours + noise estim. $60\pm 3$ $46\pm 0$ $88\pm 0$ $60\pm 2$ $57\pm 0$ $62\pm 0$
We used the MARIO OCT B-scan dataset [20] to train our architecture suitable for handling ordinal disease progression labels (Fig. 1) and first evaluated it for disease progression classification. We compared its performance to a naïve classifier trained for a 4-class problem with categorical cross-entropy. For our model, to assign each point in $\hat{y}_{\textrm{d}}$ to a class, we optimized a symmetric decision boundary around $0.5$ based on the validation data. Our model performed similar to the the naïve classifier (Table 1), with slight improvements in Rk-correlation, recall and specificity at the price of a lower F1 score mainly due to a reduced performance for the stable class. While the focus of this work was not on optimizing performance for the disease progression classification task, our model would have been placed 17/21 on the MARIO leaderboard, when evaluated on 15% of the training data in the cross-validation setting (as we did not have access to the MARIO validation set which was used to rank the participants).

### v2

@@ 3.1 Disease Progression Classification
Table 1: Classification performance (in %) for disease progression.
@@
naïve clf. $70\pm 5$ $44\pm 4$ $87\pm 1$ $60\pm 3$ $57\pm 3$ $60\pm 3$
ours $61\pm 7$ $36\pm 5$ $86\pm 1$ $59\pm 5$ $47\pm 3$ $59\pm 5$
ours + noise estim. $60\pm 7$ $36\pm 6$ $86\pm 2$ $60\pm 4$ $47\pm 4$ $60\pm 4$
We used the MARIO OCT B-scan dataset [21] to train our architecture suitable for handling ordinal disease progression labels (Fig. 1) and first evaluated it for disease progression classification on the metrics of the MARIO challenge [15]. We compared its performance to a naïve classifier (clf.) trained for a 4-class problem with categorical cross-entropy. For our model, to assign each point in $\hat{y}_{\textrm{d}}$ to a class, we optimized a symmetric decision boundary around $0.5$ based on the validation data. Our model performed comparable to the the naïve classifier (Table 1) for specificity, balanced accuracy and recall. However, in other metrics like the F1 score, our model showed reduced performance, mainly due to the performance for the stable class. While the focus of this work was not on optimizing performance for the disease progression classification task, our model would have been placed 17/21 on the MARIO leaderboard, when evaluated on 15% of the training data in the cross-validation setting (as we did not have access to the MARIO validation set which was used to rank the participants).